A rigid tank with a total volume of 0.8m3 initially contains water in the form of a saturated liquid-vapour mixture at P1=500kPa. The liquid portion of the mixture initially occupies a volume of only 0.5L (the remainder is vapour). Steam from a supply line at Pi=1000kPa and Ti=500°C is then injected into the tank until the pressure in the tank reaches P2=800kPa. At this point (final state), the mass in the tank is m2=2.8kg.

a) Find the initial quality x1 and the mass added to the tank mi during the process. (7 marks)

b) Determine the final temperature T2 in the tank. (3 marks)

c) Calculate the heat transfer during the process. (8 marks)

Tank Supply Line
Answer #1

solution? Given Data? V=0.8m3. ca) at state 1: P = 500 kPa f = 0.5L 20.0005m3. So; Vg = V-ve. =0.8-0.0005 vg D.7995m3 from th2 Mots of steam; mg spy xVg = 2.669x0.7995 -2.1338655 kg Initial Quality; x = mg mgtme - 2.1338655 2.133865570.4574565 x=0.8213 (b) from steam table at P=1000kpa and T=500°c hi=3478.3kJ/kg from steam Table at a=500kpa hf=640.1 kJ/kg hfg=2107. URJkg hmin. u (c) Acdeling to the ist law of Thermodyna Q=sutu. since there is no wok output; so wro. 80; Q=AU = maugomu. from steam